# largesymbols"0D largesymbols"0D

### v3

@@ Abstract
We explore pretraining estimators for structural econometric models. The estimator is “pretrained” in the sense that the bulk of the computational cost and researcher effort occur during its construction. Subsequent applications of the estimator to different datasets require little computational cost or researcher effort. The estimation leverages a neural net to recognize the structural model’s parameter from data patterns. As an initial trial, this paper builds a pretrained estimator for a sequential search model that is known to be difficult to estimate. We evaluate it on 12 real datasets. The estimation takes seconds to run and shows high accuracy. We provide the estimator at pnnehome.github.io. More generally, pretrained, off-the-shelf estimators can make structural models more accessible to researchers and practitioners.
@@ 1 Introduction
Pre-trained machine learning models, ranging from popular image recognition models to large language models, are achieving significant success. These pretrained models are particularly good at pattern-driven tasks that are otherwise laborious for humans. By leveraging pretrained models, end users can greatly reduce the costs of these tasks. Take image recognition for example: instead of collecting their own images, labeling them, and training a neural net from scratch, users can take a pretrained neural net and directly apply it to recognize objects in their images.
@@
This paper explores bringing pretraining to structural estimation. We develop an approach to pretrain estimators for structural econometric models. The approach builds on Wei and Jiang (2024) (more discussion later). The estimator is “pretrained” in the sense that the bulk of computational cost and researcher effort occur during the construction of the estimator. Subsequent applications of it to different datasets require little cost or effort. A pretrained estimator assumes a given structural model. Then, it uses a neural net to “recognize” the structural model’s parameter from data patterns (similar to how neural nets recognize objects in images). Therefore, we refer to it as the pretrained neural net estimator (hereafter, pretrained NNE).
Pretrained NNE makes structural modeling more accessible. Estimating a structural model is often technically complex and time-consuming. Yet, the values of structural models – such as insights on micro-founded consumer preferences and counterfactual analyses – come only after the estimation. Pretrained NNE provides an off-the-shelf tool to easily estimate a structural model. It benefits researchers and practitioners who seek to apply a structural model without delving into the technical complexities of estimation. The ease of computation can also facilitate new applications, such as integrating structural models in real-time algorithms (e.g., dynamic pricing, recommenders, bandit experiments). Pretrained estimators can be called continuously to process real-time data and feed into algorithms to adjust prices, rankings, recommendations, and more.
@@
As an initial trial, this paper focuses on sequential search models that are known to be difficult to estimate. These models predict which products a consumer will search and buy (e.g., searching products $j=1,2,8$ and buying $j=2$). One example of their applications is to model online shopping behaviors. Search models are commonly estimated by simulated maximum likelihood (SMLE), but the estimation is challenging. The vast number of search and purchase combinations makes direct likelihood simulation computationally infeasible. Workarounds have been proposed, such as likelihood smoothing or specialized simulation algorithms (see Ursu et al. 2023). Nevertheless, the computational cost of a single estimation run is still high. Moreover, multiple estimation runs are required to tune optimization routines, try starting points, tune smoothing factors, etc. Additional effort is needed to derive and code the complex simulation algorithm. More broadly, many other structural models face similar challenges in estimation.
@@
Pretrained NNE offers an alternative way for researchers to estimate a structural model. It does so by learning a direct mapping from data patterns to the structural model’s parameter. Conceptually, we can think of a pretrained NNE as follows. Many papers of structural estimation start with reduced-form evidence before estimating the structural model. Although reduced-form results do not directly translate into an estimate of the structural model’s parameter, they offer useful clues. For example, if customers are more likely to search lower-price items, we expect the price coefficient in the search model to be more negative. If customers make more searches over weekends, we know weekend customers have either a low search cost or a low outside utility. More generally, we can imagine that with a rich set of such reduced-form patterns, an expert familiar with the search model can “guesstimate” the search model parameter value. The present paper’s exercise is to leverage machine learning to create such an “expert.”
@@
We describe a standard sequential search model in Section 2. Then, in Section 3 we describe how to pretrain a NNE for this search model. At a high level, the training is based on a large number of datasets simulated from the search model. Let $\boldsymbol{\theta}$ denote the search model parameter. For any value of $\boldsymbol{\theta}$ drawn from a prior, we can use the search model to simulate a dataset ${\cal D}$. Use $\boldsymbol{m}$ to collect some data patterns of ${\cal D}$. Repeating this process $L$ times gives us $\{\boldsymbol{\theta}^{(\ell)},\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$. (We currently use $L\simeq$ 1 million). These $L$ pairs serve as the training examples for us to train a neural net mapping from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. This trained neural net is the aforementioned “expert” capable of estimating the search model. Figure 1 provides an illustration of the neural net.
We provide the pretrained NNE at: pnnehome.github.io. To apply it to estimate the search model, users only need to supply data. No further tuning is required, and the computation typically takes seconds on a laptop. Specifically, users supply a dataset ${\cal D}$ consisting of two components: (i) $\boldsymbol{y}$, which includes search and purchase choices, and (ii) $\boldsymbol{x}$, which includes product attributes that affect utility, consumer attributes that affect outside utility, and advertising attributes that affect search cost. The current version of our pretrained NNE can accommodate datasets with $n\geq 1000$ consumers, 2 to 8 product attributes, 0 to 5 consumer attributes, and 0 to 2 marketing attributes. These ranges are configurable before pretraining.
Section 4 evaluates the pretrained NNE using 12 real datasets. We conduct two sets of tests. The first set consists of Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices under a “true” $\boldsymbol{\theta}$. Pretrained NNE is faster than SMLE by 3-4 orders of magnitude. Further, it recovers $\boldsymbol{\theta}$ with accuracy equal to or higher than SMLE. The result holds even when compared against the SMLE that uses a specialized simulation algorithm for search models that takes lots of effort to implement (Jiang et al. 2021). The second set of tests estimates the search model entirely from real data. Because the true $\boldsymbol{\theta}$ is unknown, we compare the estimates by pretrained NNE with those by SMLE. Their estimates are similar. The pretrained NNE is again faster by 3-4 orders of magnitude.
Compared to the general practice of training image or language models, the pretraining of NNE exploits the specific context of structural estimation in two key aspects. First, image or language models typically must learn representations of visual or textual data to perform tasks. In contrast, we use economic theory and intuition to directly specify a data representation, $\boldsymbol{m}$, for estimating $\boldsymbol{\theta}$. For the search model, we let $\boldsymbol{m}$ include the coefficients from a set of regressions describing consumer searches and purchases. This greatly reduces the training complexity. Second, image or language models are trained with labeled images or written texts. In contrast, we generate training examples with the structural model. Specifically, we use the search model to simulate outcomes $\boldsymbol{y}$ given attributes $\boldsymbol{x}$, and we simulate $\boldsymbol{x}$ to reflect the attributes commonly seen in real applications. In particular, we use a broad range of specifications for $\boldsymbol{x}$ that vary in marginal distributions, correlations, and the sizes of within- and cross-consumer variations.
This paper extends the standard (non-pretrained) NNE in Wei and Jiang (2024). In the standard NNE, the training examples assume the observed attributes $\boldsymbol{x}$ from a given real dataset. Thus, the trained NNE is applicable specifically to that real dataset. A different real dataset requires training a separate NNE. In contrast, a pretrained NNE can be directly applied to different real datasets. [^1] Another extension to Wei and Jiang (2024) is that the current paper expands $\boldsymbol{m}$ to include not only data moments but also coefficients of reduced-form regressions. The expansion substantially improves estimation accuracy. Here, it is important to note that like the standard NNE, a pretrained NNE still assumes a specific structural model. Because the cost of building a pretrained NNE is considerably higher than training a standard NNE, pretrained estimators are perhaps best suited for structural models that are already well established in literature.
@@
From a theoretical perspective, the estimate of NNE reflects a Bayesian posterior. Building on the results in Wei and Jiang (2024), one can show that as $L\rightarrow\infty$, the pretrained NNE gives the posterior of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. This posterior is one’s best estimate of the structural model parameter after seeing (only) the data patterns. It is also known as a “limited-information” posterior, because it conditions on a set of data patterns instead of full data.
@@
Finally, we emphasize that the intention of this paper is to explore the feasibility of pretrained estimators for structural models, rather than to develop an all-around tool for estimating search models. So for the latter task, there remains room for improvement. For example, how to incorporate richer unobserved heterogeneity across consumers? How to allow instrumental variables to address possible endogeneity? Section 5 discusses these potential extensions.
@@ 1.1 Literature
This paper is related to several streams of research. First, this paper joins a recent literature that seeks to leverage machine learning for econometric estimation. In particular, the literature has studied the use of flexible machine learning models to better capture individual-level heterogeneity (e.g., Wager and Athey 2018, Athey et al. 2019, Farrell et al. 2021b), to better incorporate high-dimensional nuisance parameters (e.g., Chernozhukov et al. 2018), and to construct instrumental variables (e.g., Hartford et al. 2017, Lewis and Syrgkanis 2018, Singh et al. 2020). A main theme of this literature so far is to use machine learning to flexibly describe the relations between variables in data. The current paper differs in that the relations between data variables are described by an existing econometric model (e.g., sequential search), and we leverage machine learning to estimate the parameter of this econometric model. Recent works sharing this theme include Wei and Jiang (2024) and Kaji et al. (2023). In particular, Kaji et al. (2023) takes a substantially different approach. They borrow ideas from adversarial learning, and propose training a discriminator to construct the objective function for estimating the econometric model.
Second, we build on the successful idea of pretraining in machine learning practice. Pretraining involves training a machine learning model to learn general capabilities which can then be applied to various specific tasks. In natural language processing, pretrained models like BERT and GPT demonstrate remarkable capabilities in many language tasks (Devlin et al. 2018, Brown et al. 2020). In computer vision, pretrained neural nets such as AlexNet and ResNet have been applied widely (Krizhevsky et al. 2012, He et al. 2016). A review can be found in Han et al. (2021). Overall, pretrained models have considerably improved the productivity of researchers and practitioners. The current paper aims to bring the benefits of pretraining to structural estimation in economics and marketing. Pretrained estimators are “Train Once, Deploy Anywhere.” Their availability will make structural estimation more accessible to researchers and practitioners.
Third, this paper adds to the literature that seeks to enable model estimation while preserving data privacy. Data privacy becomes an increasingly important issue to marketers. Various methods have been proposed to preserve privacy in estimation, such as noise infusion (Abowd et al. 2012, Avella-Medina 2021), generating synthetic data (Schneider et al. 2018, Anand and Lee 2023), and estimation using aggregate data (e.g., Musalem et al. 2009). The present paper advances privacy-preserving estimation using aggregate data. Pretrained NNE uses only reduced-form data patterns to estimate a structural model. Sharing reduced-form data patterns, which are aggregated across individuals, poses far less privacy concern for data owners than sharing individual-level data.
@@ 2 Consumer search model
This section describes a standard (sequential) search model, which will be the focus of this paper. We also describe the typical data used to estimate the search model.
@@ 2.1 The model
Sequential search models describe how consumers search and choose from a set of available options. One example of the context where these models are increasingly applied is online browsing and shopping. In this context, consumer is given a webpage with a list of products; she decides which products to search (i.e., click) and which product to buy. A survey of search models can be found in Ursu et al. (2023).
@@
There are two shocks, $\mu_{ij}$ and $\varepsilon_{ij}$. The consumer observes $\mu_{ij}$ pre-search but observes $\varepsilon_{ij}$ only post-search. Both shocks are distributed as $N(0,1)$. While normalizing one shock is without loss of generality, normalizing both shocks is not. Nevertheless, many applications have normalized both shocks because it can be empirically difficult to identify both variances. We follow this practice. See Ursu et al. (2023) for more discussion.
Let vector $\boldsymbol{x}_{i}^{(cons)}$ collect the attributes of consumer $i$ (e.g., demographics, purchase history). These attributes affect the utility of an outside option as follows, with $\varepsilon_{i0}\sim N(0,1)$.
@@
Each consumer has one free search. Afterwards, a search cost is incurred for each additional search. We allow some product-level attributes that affect search cost but not utility, collected in $\boldsymbol{x}_{ij}^{(ads)}$. We will refer to them as advertising or marketing attributes. One example of such attributes is the product rankings on the webpage. The search cost $c_{ij}$ is specified as
@@
The search model predicts: (i) which products the consumer will search, (ii) which products the consumer will buy, and (iii) the order by which the consumer searches the products. We shall not delve into the details of deriving these predictions. It suffices to mention that the predictions are usually calculated using the optimal decision rule in Weitzman (1979).
It is useful to note that the search model can also be applied to contexts outside online browsing and purchase. For example, increasingly available geo-location data allow us to study consumers’ search behaviors in offline shopping. Shop visits and subsequent purchases can be modeled by a sequential search model. An interesting aspect is that the distances from shops to elevators or entrances may affect search costs (and thus be considered as a marketing attribute).

### v4

@@ Abstract
We develop pre-trained estimators for structural econometric models. The estimator uses a neural net to recognize the structural model’s parameter from data patterns. Once trained, the estimator can be shared and applied to different datasets at negligible cost and effort. Under sufficient training, the estimator converges to the Bayesian posterior given the data patterns. As an illustration, we construct a pretrained estimator for a sequential search model (available at pnnehome.github.io). Estimation takes only seconds and achieves high accuracy on 12 real datasets. More broadly, pretrained estimators can make structural models much easier to use and more accessible.
@@ 1 Introduction
Pre-trained machine learning models, ranging from popular image recognition models to large language models, are achieving significant success (e.g., AlexNet, GPT). These pretrained models allow end users to greatly reduce the costs of pattern-recognition tasks. Take image recognition for example: instead of collecting images, labeling them, and training a neural net from scratch, users can directly apply a pretrained neural net to recognize objects in their images.
@@
This paper explores the use of pretraining in structural estimation. We develop pretrained estimators for structural econometric models. The bulk of computational cost and researcher effort occur only once during the estimator’s construction. Subsequently, the estimator can be shared and applied to different datasets with negligible cost and effort. More specifically, a pretrained estimator assumes a given structural model. We use the structural model to simulate datasets, which are then used to train a neural net to recognize the structural model’s parameter from a set of data patterns (similar to how neural nets recognize objects from images). With sufficient training, the estimator converges to the Bayesian posterior of the parameter given the data patterns. We refer to the estimator as the pretrained neural net estimator (hereafter, pretrained NNE).
Pretrained NNE makes structural models easier to use and more accessible. In practice, estimating a structural model is often technically complex and time-consuming. Yet, the value of structural models – such as uncovering micro-founded preferences and enabling counterfactuals – is realized only after estimation. A pretrained NNE can be built once and then shared, allowing its users to easily and quickly estimate a structural model across applications. For users, the cost and effort of estimation become comparable to those in reduced-form regressions. Further, pretrained estimators can broaden the scope of applications of structural models. For example, one may apply pretrained estimators continuously to real-time data, enabling the integration of structural models into real-time algorithms (e.g., dynamic pricing, recommenders, bandit experiments). Another benefit of pretrained NNE is that it helps preserve privacy in estimation because it relies only on aggregate data patterns as input.
@@
As an illustration, this paper focuses on sequential search models. These models characterize consumer search and purchase choices, and are finding growing application especially with online shopping data. Search models are estimated by simulated maximum likelihood (SMLE), but estimation is challenging. The vast number of search and purchase combinations makes it infeasible to directly simulate likelihood. To this end, the literature has introduced several workarounds over time, such as likelihood smoothing or specialized simulation algorithms (see Ursu et al. 2023). Even then, a single estimation run is expensive, let alone the multiple runs needed to tune optimizers, try starting points, tune smoothing factors, tune simulation algorithms, etc. Importantly, these tunings have to be performed separately for each application. More broadly, similar estimation challenges exist in many other structural models.
@@
Pretrained NNE offers an alternative way to estimate a structural model. It does so by learning a direct mapping from data patterns to the structural model’s parameter. Conceptually, we can think of a pretrained NNE as follows. Research papers using structural estimation often start with reduced-form evidence before estimating the structural model. Although reduced-form results do not directly translate into an estimate of the structural model’s parameter, they offer useful clues. For example, if customers are more likely to search lower-price items, we expect the price coefficient in the search model to be more negative. If customers make more searches over weekends, we know weekend customers have either a low search cost or a low outside utility. More generally, we can imagine that with a rich set of such reduced-form patterns, an expert familiar with the search model can guesstimate the search model parameter value. The present paper’s exercise is to leverage machine learning to create such an “expert.”
@@
We describe a standard sequential search model in Section 2. Then, in Section 3 we describe how to pretrain an NNE for this search model. At a high level, the training uses a large number of datasets simulated from the search model. Let $\boldsymbol{\theta}$ denote the search model parameter. For each draw of $\boldsymbol{\theta}$ from a prior, we can use the search model to simulate a dataset $\cal D$}. Use $\boldsymbol{m}$ to collect some data patterns of $\cal D$} (e.g., summary statistics, regression coefficients). Repeating this process for $L$ draws of $\boldsymbol{\theta}$ gives us $\{\boldsymbol{\theta}^{(\ell)},\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$ (we currently use $L\simeq$ 1 million). These $L$ pairs serve as the training examples for a neural net to learn a mapping from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. The trained neural net becomes the aforementioned “expert” capable of estimating the search model. Figure 1 offers an illustration of this neural net.
@@
From a theoretical perspective, the estimate of NNE reflects a Bayesian posterior. Under mild conditions, as $L\rightarrow\infty$, the pretrained NNE converges to the posterior mean of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. This mean is known as a limited-information posterior mean, and it is one’s best estimate for the structural model’s parameter after seeing the data patterns.
We provide our pretrained NNE at: pnnehome.github.io. It can be applied off-the-shelf to estimate the search model: users only need to supply data, no further tuning is required, and the computation typically takes seconds on a laptop. Specifically, users supply a dataset $\cal D$} consisting of two components: (i) $\boldsymbol{y}$, which includes search and purchase choices, and (ii) $\boldsymbol{x}$, which includes product attributes that affect utility, consumer attributes that affect outside utility, and advertising attributes that affect search cost. The current version of our pretrained NNE can accommodate datasets with $n\geq 1000$ consumers, 2 to 8 product attributes, 0 to 5 consumer attributes, and 0 to 2 advertising attributes. These ranges are configurable before pretraining.
Section 4 evaluates the pretrained NNE using 12 real datasets. We conduct two sets of tests. The first set consists of Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices under a “true” $\boldsymbol{\theta}$. Pretrained NNE is faster than SMLE by 3-4 orders of magnitude. Further, it recovers $\boldsymbol{\theta}$ with accuracy equal to or higher than SMLE. The result holds even when compared against the SMLE that uses a specialized simulation algorithm for search models that takes considerable effort to implement (Jiang et al. 2021). The second set of tests estimates the search model entirely from real data. Because the true $\boldsymbol{\theta}$ is unknown, we compare the pretrained NNE’s estimates with SMLE’s estimates. Their estimates turn out to be similar. The pretrained NNE is again faster by 3-4 orders of magnitude.
An important detail in pretraining NNE is to specify the data patterns $\boldsymbol{m}$ — a representation of the input $\cal D$}. In computer vision or language processing, representations of inputs (images or texts) typically must be learned via training. By contrast, in structural models, economic theory and intuition often provide sufficient guidance for us to specify an $\boldsymbol{m}$ that can identify $\boldsymbol{\theta}$. This greatly reduces the training complexity. To optimize estimation efficiency, alternative specifications of $\boldsymbol{m}$ can be compared using the neural net’s accuracy on validation examples (i.e., a held-out subset of training examples). Importantly, this effort to select $\boldsymbol{m}$ arises only in pretraining and is not borne by the users of a pretrained NNE. For our search model, we choose $\boldsymbol{m}$ to include the coefficients from a set of regressions that describe consumer search and purchase behaviors. [^1] In Wei and Jiang (2024), $\boldsymbol{m}$ includes only data moments. Expanding $\boldsymbol{m}$ to include reduced-form regression coefficients leads to a substantial improvement in estimation accuracy for the search model.
This paper extends the standard (non-pretrained) NNE in Wei and Jiang (2024). In a standard NNE, all training examples assume the same observed attributes $\boldsymbol{x}$ from a real dataset. Thus, the trained net is applicable only to that dataset. In contrast, a pretrained NNE is applicable across different datasets. To achieve this, in each training example we draw $\boldsymbol{x}$ from a distribution (and then use the search model to simulate $\boldsymbol{y}$ given the drawn $\boldsymbol{x}$). This distribution serves the role of a prior for $\boldsymbol{x}$. We specify it to reflect attributes common in real applications, covering diverse scenarios that vary in the number of attributes, marginal distributions, correlations, and within- versus cross-consumer variation. We use a large $L$ to cover these scenarios well. It is important to note that an NNE (standard or pretrained) still assumes a specific structural model. Therefore, pretrained NNE is more suited for established structural models that are poised for broader applications, whereas standard NNE is more suited for newly proposed or ad hoc structural models.
@@
Finally, we emphasize that this paper’s goal is to explore the feasibility of pretrained estimators for structural models, rather than to develop a full-fledged tool for estimating search models. For the latter task, there remains room for improvement, such as allowing richer unobserved consumer heterogeneity and incorporating instruments to address possible endogeneity. These extensions are non-trivial and we discuss them in Section 5.
@@ 1.1 Literature
This paper is related to several streams of research. First, this paper joins a recent literature that seeks to leverage machine learning for econometric estimation. In particular, the literature has studied the use of flexible machine learning models to better capture individual-level heterogeneity (e.g., Wager and Athey 2018, Athey et al. 2019, Farrell et al. 2021b), to better incorporate high-dimensional nuisance parameters (e.g., Chernozhukov et al. 2018), and to construct instrumental variables (e.g., Hartford et al. 2017, Lewis and Syrgkanis 2018, Singh et al. 2020). A main theme of this literature so far is to use machine learning to flexibly describe the relations between variables in data. The current paper differs in that the relations between data variables are described by an existing econometric model (e.g., sequential search), and we leverage machine learning to estimate the parameter of this econometric model. Recent works sharing this theme include Wei and Jiang (2024) and Kaji et al. (2023). In particular, Kaji et al. (2023) borrow ideas from adversarial learning, and propose training a discriminator to construct the objective function for estimating the econometric model. This approach is conceptually different from ours.
Second, we build on the successful idea of pretraining in machine learning practice. Pretraining involves training a machine learning model to learn general capabilities which can then be applied to various specific tasks. In natural language processing, pretrained models like BERT and GPT demonstrate remarkable capabilities in many language tasks (Devlin et al. 2018, Brown et al. 2020). In computer vision, pretrained neural nets such as AlexNet and ResNet have been applied widely (Krizhevsky et al. 2012, He et al. 2016). A review can be found in Han et al. (2021). Pretrained models have considerably improved the productivity of researchers and practitioners. The current paper aims to bring the benefits of pretraining to structural estimation in economics and marketing. Pretrained estimators are “Train Once, Deploy Anywhere.” Their availability will make structural estimation more accessible to researchers and practitioners.
Third, this paper adds to the literature that seeks to enable model estimation while preserving data privacy. Data privacy is becoming increasingly relevant to marketers. Various methods have been proposed to preserve privacy in estimation, such as noise infusion (Abowd et al. 2012, Avella-Medina 2021), generating synthetic data (Schneider et al. 2018, Anand and Lee 2023), and estimation using aggregate data (e.g., Musalem et al. 2009). The present paper advances privacy-preserving estimation using aggregate data. Pretrained NNE requires only reduced-form data patterns to estimate a structural model. Sharing reduced-form data patterns, which are aggregated across individuals, poses far less privacy concern for data owners than sharing individual-level data.
@@ 2 Consumer search model
In this section, we describe a standard (sequential) search model and the typical data used to estimate it. The next section will present the pretrained NNE for this search model.
@@ 2.1 The model
Sequential search models describe how consumers search and choose from a set of available options. One example where these models are increasingly applied is online browsing and shopping. In this context, consumer is given a webpage with a list of products; she decides which products to search (i.e., click) and which product to buy. A survey of sequential search models can be found in Ursu et al. (2023).
@@
There are two shocks, $\mu_{ij}$ and $\varepsilon_{ij}$. The consumer observes $\mu_{ij}$ pre-search but observes $\varepsilon_{ij}$ only post-search. Both shocks are distributed as $N(0,1)$. While normalizing one shock is without loss of generality, normalizing both shocks is not. Nevertheless, many applications have normalized both shocks because it can be empirically difficult to identify the variances of both shocks. We follow this practice. See Ursu et al. (2023) for more discussion.
Let vector $\boldsymbol{x}_{i}^{(cons)}$ collect the attributes of consumer $i$ (e.g., demographics, purchase history). These attributes affect the utility of an outside option as follows. The shock $\varepsilon_{i0}$ is known pre-search and follows $N(0,1)$.
@@
Each consumer has one free search. Afterwards, a search cost is incurred for each additional search. We allow for some product-level attributes that affect search cost but not utility, collected in $\boldsymbol{x}_{ij}^{(ads)}$. We will refer to them as advertising or marketing attributes. One example of such attributes is the product rankings on the webpage. The search cost $c_{ij}$ is specified as
@@
The search model characterizes: (i) which products the consumer will search, (ii) which products the consumer will buy, and (iii) the order in which the consumer searches the products. We shall not delve into deriving the optimal choices of the consumer. It suffices to note that these choices are typically computed using the optimal decision rule in Weitzman (1979).
It is useful to note that the search model can also be applied to contexts outside online browsing and purchase. For example, geo-location data are increasingly available. They allow one to study consumers’ search behaviors in offline shopping. Shop visits and subsequent purchases can be modeled by a sequential search model. An interesting aspect is that the distances from shops to elevators or entrances may affect search costs (and thus be considered as a marketing attribute).